The joint probability density of X, Y is

f(x, y) = e −(x + y) xi > 0 i = 1, 2

0 otherwise

Using the change of variable technique, determine the joint distribution of Z = X and W = X + Y


1
Expert's answer
2022-05-12T08:56:55-0400

At first, we check that "f(x,y)" is a valid joint probability density. We have to prove that "\int_0^{+\infty}\int_0^{+\infty}e^{-(x+y)}dxdy=1". Compute the integral at the left side: "\int_0^{+\infty}\int_0^{+\infty}e^{-(x+y)}dxdy=\int_{0}^{+\infty}-e^{-(x+y)}|_0^{+\
\infty}dy=\int_{0}^{+\infty}e^{-y}dy=-e^{-y}|_0^{+\infty}=1".

Thus, it is the valid distribution. In order to find the joint distribution, consider the probability: "P\{X<z,X+Y<w\}". In case "z<0" or "w<0", "P\{X<z,X+Y<w\}=0". In case "z>0,w>0", we get: "P\{Z<z,X+Y<w\}=\iint_{x<z,x+y<w}f(x,y)dxdy=\int_0^{min(z,w)}(\int_0^{w-x}e^{-(x+y)}dy)dx=\int_0^{min(z,w)}-e^{-(x+y)}|_0^{w-x}dx=\int_0^{min(z,w)}(e^{-x}-e^{-w})dx=-e^{-x}|_0^{min(z,w)}-min(z,w)e^{-w}=1-e^{-min(z,w)}-min(z,w)e^{-w}."

Thus, in case "z<w", we get: "P\{Z<z,X+Y<w\}=1-e^{-z}-ze^{-w}". For "z>w" we have: "P\{Z<z,X+Y<w\}=1-e^{-w}-we^{-w}". Denote "F(z,w)=P\{Z<z,X+Y<w\}"

The joint probability density is: "p(z,w)=\frac{\partial^2F(z,w)}{\partial z\partial w}=e^{-w}", "z<w", "z>0,w>0".

Answer: The joint probability density function is: "p(z,w)=e^{-w}", "z<w", "z>0,w>0."